# TIDMAD: Time Series Dataset for Discovering Dark Matter with AI Denoising

### v1

@@ 1 Introduction
Due to its extremely rare interactions with normal matter, the signal of dark matter is often submerged in a sea of noise from various sources internal and external to the experimental apparatus. Machine learning (ML) offers a promising means to reduce this noise. Advancements in denoising techniques using ML algorithms have the potential to significantly improve dark matter analyses [12]. These techniques enable the detection of weaker dark matter signals, or in the case of no observation, the setting of stronger limits. In other words, improvements in data denoising directly enhance the scientific reach of dark matter experiments. In this paper, we present an ultra-long time series dataset produced by a real dark matter detector: ABRACADABRA (A Broadband/Resonant Approach to Cosmic Axion Detection with an Amplifying B-field Ring Apparatus, abbrev. ABRA). ABRA is the world leading sub-$\mu eV$ dark matter experiment that pioneered the quantum enabled lumped element dark matter detection technique [8, 9, 13]. We operated the ABRA detector in February 2024 to obtain a special time series dataset for these studies: TIDMAD (TIme series dataset for discovering Dark Matter with Ai Denoising). It can be partitioned into three parts: (1) training data, (2) validation data, and (3) science data.
The training data include time series data where a dark matter-like signal is injected by hardware. If dark matter enters ABRA, it will manifest itself as a sinusoidal oscillation mode within the time series; therefore, the injected signals are also sinusoidal oscillations at specific frequencies. Both the detected (noisy) time series and the injected (ground truth, clean) time series are provided with one-to-one temporal correspondence. This allows the training of machine learning algorithms to denoise the detector data and recover the injected signal. The validation data is used to produce Benchmark 1: Denoising Score, see Section 4.1. Algorithms that effectively dampen the detector noise while amplifying the injected signal will achieve a better denoising score. The science data is collected without the injected signal with an extended duration to produce Benchmark 2: Dark Matter Limits, as discussed in the previous paragraph. The limit generation procedure is detailed in Section 4.2. Several traditional and deep learning denoising algorithms are presented in Section 3 and Appendix B, where the resulting denoised data is benchmarked against the raw, un-denoised detector data.
@@ 1.1 Axion dark matter and ABRACADABRA
In recent years, the axion has emerged as one of the leading dark matter candidates due to its theoretical elegance. Axions interact with normal matter via electromagnetism, which can be characterized by a physics parameter $g_{a\gamma\gamma}$. Due to its small mass $m_{a}<1eV$ ($10^{-6}$ times smaller than electron), axions acts as a classical field oscillating at a frequency $f_{a}=m_{a}/2\pi$. Astrophysical measurements determine that the Earth exists in a bath of dark matter with a known local density of $\rho_{DM}$ [4].
Figure 1: Left to right: ABRA dilution fridge with outer vacuum cans on; Coldest stage of ABRA fridge above shielded 1T superconducting torroidal magnet; Interior of ABRA magnet including pickup and calibration loop wires existing the center of the magnet; Effective circuit diagram for both dark matter and injected signals.
The latest advancements in quantum detector technology have facilitated new avenues to search for the axion. ABRA is one of the novel detectors designed to search for axions leveraging these advancements in quantum technologies [9]. ABRA capitalizes on the fact that we are immersed in a bath of a feebly electromagnetically-interacting, oscillating dark matter field to detect this elusive particle. Specifically, in the presences of a static magnetic field $\textbf{B}_{0}$, the axion, henceforth referred to as dark matter, induces an oscillating magnetic field $\textbf{B}_{a}$. Thus, to detect dark matter ABRA provides a strong magnetic field $B_{0}=1T$ and uses a superconducting pickup loop to read out the oscillating dark matter signal. Read out by a superconducting quantum interference device (SQUID), the pickup loop detects the dark matter signal as a time-oscillating current given by
@@ 1.2 TIDMAD construction
In the classical analysis, we use a calibration procedure to determine the end-to-end response of our system for different signal frequencies. As shown in Figure 1, the ABRA detector contains a toroidal magnet equipped with both a pickup loop and a calibration loop. During calibration, we first inject a fake dark matter signal into the calibration loop at a specific frequency. This generates a sine wave with a known amplitude and frequency, creating a dark matter-like flux with our pickup loop. Finally, this flux is detected by the SQUID sensor for detector calibration.
@@
Figure 2: 10-millisecond snapshot of the time series in TIDMAD training dataset.
@@
The TIDMAD dataset presented in this work is inspired by this calibration procedure. The ABRA detector hardware enables us to simultaneously record two types of ultra-long time series: the one injected into the calibration loop (referred to as the “injected time series”) and the one detected by the SQUID sensor coupled to the pickup loop (referred to as the “SQUID time series”). As shown in Figure 2, the injected time series exhibit a clear sinusoidal oscillatory signal, which can be considered the ground truth. Meanwhile, the SQUID time series contains the same ground truth submerged within a sea of detector noises. The two time series are exactly aligned at every time step. This defines the signal recovery task: a model could be applied to the SQUID time series to reproduce the injected signal in the injected time series. A model trained on this task will be efficient in rejecting noise of different kinds while retaining the dark-matter-like signal within the detector. We then collected a science dataset where no fake dark matter signal is injected. The trained denoising model can then be applied to the SQUID time series of the science dataset. If a sinusodial signal is found after denoising, it could potentially be a real dark matter particle entering the detector.
2 Dataset construction
@@
The data presented in this paper were acquired using the ABRACADABRA detector at MIT. The overall schematics of the data is shown in Figure 3. All data are saved as a series at $10$ MS/s (Megasample per second), where each sample is a 8-bit integer ranging from -128 to 127. These integers can be converted into a physics units of mV (millivolts) with a scaling factor of $40/128$. The procured datasets are stored at Open Science Data Federation (OSDF) in .hdf5 format and can be accessed via the download_data.py script in the github repository provided.
@@
Figure 3: TIDMAD data flow explaining how data splits (green) result in benchmarks (blue). Rectangles correspond to provided scripts. Left to right, top to bottom the scripts are train.py, inference.py, benchmark.py, process_science_data.py, brazilband.py
@@ 3 Experiments
We adopted five different denoising algorithms to the dataset to achieve signal recovery, including two traditional algorithms and three deep learning models. The traditional model can be directly applied to the validation dataset using inference.py, while deep learning models need to be trained first on the training dataset with train.py. The five algorithms are listed below:
  • Moving average: a simple moving average with a window size of 100, implemented using the numpy convolve function.
  • Savitzky-Golay filter: with a window size of 100 and a polynomial order of 11.
@@
  • Transformer: the transformer utilizes a self-attention mechanism to capture long-distance dependencies in sequences  [16]. After processing by the multi-layer Transformer encoder, the model effectively extracts features and represents the input sequence. Finally, the encoded sequence is mapped to the output dimension through a linear layer for the same 256-class classification decision as PU-Net. Positional encoding is also added before the time series is fed into Transformer layers. Both PU-Net and Transformer are trained using Focal Loss to handle class-imbalanced segmentation labels  [6].
The benchmarking results of these models are discussed in Section 4. There are two additional constraints for the deep learning models. First, due to memory constraints, we segment the ultra-long SQUID and injected time series into smaller segments before feeding them into each model. The exact segment sizes are outlined in Table 1. Secondly, due to the large range of injected frequencies, training multiple versions of the same model to handle different frequency ranges is necessary. Both limitations and additional details of the deep learning models are discussed in Appendix  A.
@@ 4.1 Benchmark 1: denoising score
The denoising score is a modified signal-to-noise ratio (SNR) of the denoised CH1 SQUID time series. It is calculated over the validation dataset by first segmenting both the injected and SQUID time series into one-second segments. Each second of the time series is transformed into a power spectral density (PSD) using a squared fast Fourier transform. This frequency domain data records signal power at each frequency – $PSD(\nu)$. Since the injected dark matter signal is a clean sinusodial oscillation, it should appear as a single-bin peak ($\nu_{0}$) in the PSD, while noise in SQUID time series is distributed across all frequency bins. The location of $\nu_{0}$ is identified by the PSD of the injected time series (ground truth) as the largest single bin peak relative to its nearest neighbors.
[…]
@@
We examine the validity of $\Lambda$ to represent denoising efficiency through a study involving added Gaussian noise. In this study, Gaussian noise is imposed on the injected time series, and $\Lambda$ are calculated at different levels of Gaussian noise amplitudes and standard deviations. As shown in Figure 4 (right), $\Lambda$ exhibits an exponential decay trend with increased Gaussian noise amplitude. To establish a linear correlation between the denoising score and noise, we apply a logarithmic transformation to $\Lambda$ to calculate the final denoising score:
Denoising $\text{Denoising Score}=log_{5.27}\Lambda$ (7)
@@
Table 1: Fine and coarse denoising score for raw data, traditional algorithms, and trained ML models
Algorithms Segment Size Parameters Fine Score Coarse Score
@@
Moving Average $1\times 10^{6}$ window = 100 0.52 0.64
SG Filter $1\times 10^{6}$ window = 100, order = 11 -2.77 -2.35
FC Net $4\times 10^{4}$ See Appendix  A 6.43 6.55
PU Net $4\times 10^{4}$ See Appendix  A 3.69 3.84
Transformer $2\times 10^{4}$ See Appendix  A 3.95 4.18
Table 1 shows the denoising scores for all algorithms discussed in Section 3. The case with no denoising is shown in the first row, with its fine denoising score calibrated to 1. We observed that in all cases, the coarse denoising score is slightly higher than the fine denoising score. Based on the results, all traditional algorithms decrease the denoising score as time domain averaging erases high-frequency signals in the region of interest. Meanwhile, all deep learning algorithms efficiently boost the denoising score. Surprisingly, we observed that the FC Net model achieved the best performance with a denoising score of 6.43.
@@ 4.2 Benchmark 2: dark matter limit
The second benchmark empowers algorithm creators with the capability to conduct a community-standard dark matter search using the science dataset. Following the discussion in the Introduction, the two physics parameters for the dark matter candidate in this paper are the dark matter mass ($m_{a}$) and the dark matter to electromagnetic coupling ($g_{a\gamma\gamma}$). Null results from different dark matter experiments place limits within this parameter space expressed by the shaded regions in Figure 5. In the physics community, a better dark matter limit is represented by pushing towards lower values of $g_{a\gamma\gamma}$ at different $m_{a}$. We provide a comprehensive tool necessary for performing the statistical analysis to produce dark matter limits in Figure 5. The dark matter mass $m_{a}$ is directly proportional to frequency, therefore the limit-setting procedure is repeated for 11.1 million independent $m_{a}$ from 0.4 neV (100 kHz) to 8 neV (2 MHz). The dark matter limit at each $m_{a}$ is obtained using a frequentist log-likelihood ratio test statistic (TS), with the results depicted as regions in Figure 5. The detail of this analysis can be found in Appendix  B.
This analysis is performed twice on the SQUID time series of the science dataset: once without any denoising algorithm which produces the ABRA-TIDMAD Raw limit, and once with FC Net, the top-performing denoising algorithm, which produces the ABRA-TIDMAD Denoised limit. These limits can be directly compared to the previous world-leading ABRA Run 3 limit, limits obtained by other dark matter experiments, as well as theoretical predictions [13]. While the ABRA-TIDMAD Denoised limit does not outperform the ABRA Run 3 limit due to hardware and time constraints, it is evident that denoising algorithms significantly improved the dark matter limit by 1-2 orders of magnitude across different $m_{a}$. Although the size of the ABRA-TIDMAD science dataset is only 1% of the ABRA Run 3 science dataset, the AI denoising algorithm boosted the ABRA-TIDMAD limit to nearly the same level as ABRA Run 3 and even surpassed the ABRA Run 3 limits at small $m_{a}$.
Figure 5: Plotted alongside the present, state-of-the-art axion dark matter limits (grey) are the $2\sigma$ exclusion limits for ABRA Run 3 (pink) [13], ABRA-TIDMAD Raw limit (red), and ABRA-TIDMAD Denoised limit from the trained FC net (maroon). Shaded regions correspond to pairing of dark matter model parameters ($m_{a}$, $g_{a\gamma\gamma}$) that are ruled out by the specified experiment or observation, and the bright yellow region indicates theoretical predictions. The plotting script is modified from the publicly available AxionLimits repository [7].

### v2

@@ 1 Introduction
Attributable to its extremely rare interactions with normal matter, the signal of dark matter is often submerged in a sea of noise from various sources internal and external to the experimental apparatus. Machine learning (ML) offers a promising means to reduce this noise. Advancements in denoising techniques using ML algorithms have the potential to significantly improve dark matter analyses [24]. These techniques enable the detection of weaker dark matter signals, or in the case of no observation, the setting of stronger limits. In other words, improvements in data denoising directly enhance the scientific reach of dark matter experiments. In this paper, we present an ultra-long time series dataset produced by a real dark matter detector: ABRACADABRA (A Broadband/Resonant Approach to Cosmic Axion Detection with an Amplifying B-field Ring Apparatus, abbrev. ABRA-10cm). ABRA-10cm is the world leading sub-$\mu eV$ dark matter experiment that pioneered the quantum enabled lumped element dark matter detection technique [19, 20, 25]. We operated the ABRA-10cm detector in February 2024 to obtain a special time series dataset for these studies: TIDMAD (TIme series dataset for discovering Dark Matter with Ai Denoising). It can be partitioned into three parts: (1) training data, (2) validation data, and (3) science data.
The training data include time series data where a dark matter-like signal is injected by hardware. If dark matter enters ABRA, it will manifest itself as a sinusoidal oscillation mode within the time series; therefore, the injected signals are also sinusoidal oscillations at specific frequencies. Both the detected (noisy) time series and the injected (ground truth, clean) time series are provided with one-to-one temporal correspondence. This allows the training of machine learning algorithms to denoise the detector data and recover the injected signal. The validation data is used to produce Benchmark 1: Denoising Score, see Section 4.1. Algorithms that effectively dampen the detector noise while amplifying the injected signal will achieve a better denoising score. The science data is collected without the injected signal with an extended duration to produce Benchmark 2: Dark Matter Limits, as discussed in the previous paragraph. The limit generation procedure is detailed in Section 4.2. The scientific data are titled to reflect their use in producing real, community-standard physics results that are suitable for presentation in scientific journals. Several traditional and deep learning denoising algorithms are presented in Section 3 and Appendix C, where the resulting denoised data is benchmarked against the raw, un-denoised detector data.
@@ 1.1 Axion dark matter and ABRACADABRA
In recent years, the axion has emerged as one of the leading dark matter candidates as a result of its theoretical elegance. Axions interact with normal matter via electromagnetism, which can be characterized by a physics parameter $g_{a\gamma\gamma}$. Arising from its small mass $m_{a}<1eV$ ($10^{-6}$ times smaller than electron), axions acts as a classical field oscillating at a frequency $f_{a}=m_{a}/2\pi$. Astrophysical measurements determine that the Earth exists in a bath of dark matter with a known local density of $\rho_{DM}$ [11].
Figure 1: Left to right: ABRA-10cm dilution fridge with outer vacuum cans on; Coldest stage of ABRA-10cm fridge above shielded 1T superconducting torroidal magnet; Interior of ABRA-10cm magnet including pickup and calibration loop wires existing the center of the magnet; Effective circuit diagram for both dark matter and injected signals.
The latest advancements in quantum detector technology have facilitated new avenues to search for the axion. ABRA-10cm is one of the novel detectors designed to search for axions leveraging these advancements in quantum technologies [20]. ABRA-10cm capitalizes on the fact that we are immersed in a bath of a feebly electromagnetically-interacting, oscillating dark matter field to detect this elusive particle. Specifically, in the presences of a static magnetic field $\textbf{B}_{0}$, the axion, henceforth referred to as dark matter, induces an oscillating magnetic field $\textbf{B}_{a}$. Thus, to detect dark matter ABRA-10cm provides a strong magnetic field $B_{0}=1T$ and uses a superconducting pickup loop to read out the oscillating dark matter signal. Read out by a superconducting quantum interference device (SQUID), the pickup loop detects the dark matter signal as a time-oscillating current given by
@@ 1.2 TIDMAD construction
In the classical analysis, we use a calibration procedure to determine the end-to-end response of our system for different signal frequencies. As shown in Figure 1, the ABRA-10cm detector contains a toroidal magnet equipped with both a pickup loop and a calibration loop. During calibration, we first inject a fake dark matter signal into the calibration loop at a specific frequency. This generates a sine wave with a known amplitude and frequency, creating a dark matter-like flux with our pickup loop. Finally, this flux is detected by the SQUID sensor for detector calibration.
The dark matter signal injected into the calibration loop by the signal generator follows the form prescribed by axion theory, as shown in Equation 1 and derived in Appendix A. We specifically choose to inject sine waves with frequencies from 1.1 kHz to 4.9 MHz, corresponding to axion masses $m_{a}=[0.005,17]$ neV, to target the mass range that our experimental hardware is designed to detect. The injected signals were all set to an amplitude of 50 mV to ensure a reasonable signal-to-noise ratio. A total of 309 different frequencies were sequentially stepped through, from 1.1 kHz to 4.9 MHz, simulating 309 distinct axion masses in our detector hardware.
@@
Figure 2: 10-millisecond snapshot of the time series in TIDMAD training dataset compared to the power spectral density of the same data snapshot.
@@
The TIDMAD dataset presented in this work is inspired by this calibration procedure. The ABRA-10cm detector hardware enables us to simultaneously record two types of ultra-long time series: the one injected into the calibration loop (referred to as the “injected time series”) and the one detected by the SQUID sensor coupled to the pickup loop (referred to as the “SQUID time series”). As shown in Figure 2, the injected time series exhibit a clear sinusoidal oscillatory signal, which can be considered the ground truth. Meanwhile, the SQUID time series contains the same ground truth submerged within a sea of detector noises. While the ground truth time series is simple, the detector readout contains a complex spectrum of noise spanning a large frequency range, which cannot be modeled using any simple approach [10]. The two time series are exactly aligned at every time step. This defines the signal recovery task: a model could be applied to the SQUID time series to reproduce the injected signal in the injected time series. A model trained on this task will be efficient in rejecting noise of different kinds while retaining the dark-matter-like signal within the detector. We then collected a science dataset where no fake dark matter signal is injected. The trained denoising model can then be applied to the SQUID time series of the science dataset. If a sinusoidal signal is found after denoising, it could potentially be a real dark matter particle entering the detector.
2 Dataset description
@@
The data presented in this paper were acquired using the ABRACADABRA detector. The overall schematics of the data is shown in Figure LABEL:fig:dataflow. All data are saved as a series at $10$ MS/s (Megasample per second), where each sample is a 8-bit integer ranging from -128 to 127. These integers can be converted into a physics units of mV (millivolts) with a scaling factor of $40/128$. The procured datasets are stored at Open Science Data Federation (OSDF) [34] in .hdf5 format and can be accessed via the download_data.py script in the github repository provided.
@@
Figure 3: TIDMAD data flow explaining how data splits (green) result in benchmarks (blue). Rectangles correspond to provided scripts. Left to right, top to bottom the scripts are , , , , fig:dataflow
@@ 3 Experiments
We experimented five different denoising algorithms including two traditional algorithms and three deep learning models. The traditional algorithm can be directly applied to the validation dataset using inference.py, while deep learning models need to be trained first on the training dataset with train.py. The five algorithms are listed below:
  • Moving average: a simple moving average with a window size of 19, implemented using the numpy convolve function.
  • Savitzky-Golay filter: with a window size of 19 and a polynomial order of 11
  • Fourier Averaging (baseline): This operation was originally adopted by the ABRACADABRA experiment to compress its data [19, 25]. For every 10-second time series segment, a Fourier transform is applied to each one-second interval, and the resulting 10 Fourier spectra are averaged to produce a single spectrum representing the 10-second segment. This method is considered the baseline approach, representing the current status quo of ABRACADABRA.
@@
  • WaveNet [30]: a deep neural network that was originally developed to generates high-quality raw audio waveforms by autoregressively predicting each sample using dilated causal convolutions. The model was adapted with residual blocks containing exponentially increasing dilation rates, gated activations, and skip connections to maintain the same input-output shape for TIDMAD denoising task.
@@
  • Transformer: the transformer utilizes a self-attention mechanism to capture long-distance dependencies in sequences  [31]. After processing by the multi-layer Transformer encoder, the model effectively extracts features and represents the input sequence. Finally, the encoded sequence is mapped to the output dimension through a linear layer for the same 256-class classification decision as PU-Net. Positional encoding is also added before the time series is fed into Transformer layers.
  • RNN Sequence to Sequence Model [30]: a neural architecture with separate encoder and decoder RNNs where the encoder processes the entire input sequence to produce a context representation, and the decoder autoregressively generates the output sequence. Both encoder and decoder adopts a LSTM architecture, and the decoder output is the same 256-class classification decision as PU-Net. PU-Net, Transformer, WaveNet, and RNN Seq2Seq are all trained using Focal Loss to handle class-imbalanced segmentation labels  [16].
The benchmarking results of these models are discussed in Section 4. A hyperparameter search study is conducted on the Moving Average and Savitzky-Golay filter. The best-performing window size and polynomial order are chosen, with the rest of the study detailed in Appendix E. There are two additional constraints for the deep learning models. First, because of memory constraints, we segment the ultra-long SQUID and injected time series into smaller segments before feeding them into each model. The exact segment sizes are outlined in Table 1. Secondly, due to the broad frequency spectrum in our input data, we implemented frequency splitting for all models except WaveNet, training multiple specialized versions of each model to handle distinct frequency ranges. During benchmarking, we observed that a single WaveNet model is efficient to handle all frequencies. Both limitations and additional details of the deep learning models are discussed in Appendix  B.
@@ 4.1 Benchmark 1: denoising score
The denoising score is a modified signal-to-noise ratio (SNR) of the denoised CH1 SQUID time series. It is calculated over the validation dataset by first segmenting both the injected and SQUID time series into one-second segments. Each second of the time series is transformed into a power spectral density (PSD) using a squared fast Fourier transform. This frequency domain data records signal power at each frequency -- $PSD(\nu)$. Since the injected dark matter signal is a clean sinusodial oscillation, it should appear as a single-bin peak ($\nu_{0}$) in the PSD, while noise in SQUID time series is distributed across all frequency bins. The location of $\nu_{0}$ is identified by the PSD of the injected time series (ground truth) as the largest single bin peak relative to its nearest neighbors.
$\nu_{0}=\underset{\nu}{\mathrm{argmax}}\Big(PSD_{\mathrm{Injected}}(\nu)-(PSD_{\mathrm{Injected}}(\nu-df)+PSD_{\mathrm{Injected}}(\nu+df))\Big)$ (3)
@@
We examine the validity of $\Lambda$ to represent denoising efficiency through a study involving added Gaussian noise. In the ABRA-10cm detector, there are a multitude of independent, random noise sources internal and external to the detector thus the noise in our data follows a Gaussin distribution. In this study, Gaussian noise is imposed on the injected time series, and $\Lambda$ are calculated at different levels of Gaussian noise amplitudes and standard deviations. As shown in Figure 4 (right), $\Lambda$ exhibits an exponential decay trend with increased Gaussian noise amplitude. To establish a linear correlation between the denoising score and noise, we apply a logarithmic transformation to $\Lambda$ to calculate the final denoising score:
$\text{Denoising Score}=log_{5.27}\Lambda$ (7)
@@
Table 1: Fine and coarse denoising score for raw data, traditional algorithms, and trained ML models. FS means frequency splitting, or training multiple versions of the same model to handle different frequency ranges. The detail of segment size and FS is discussed in Section B.
Algorithms Segment Size FS Parameters Fine Score Coarse Score
@@
Fourier Averaging $1\times 10^{8}$ N/A 10-fold Average 0.24 0.26
Moving Average $1\times 10^{6}$ N/A window = 100 0.86 0.95
SG Filter $1\times 10^{6}$ N/A window = 19, order = 11 0.95 1.04
FC Net $4\times 10^{4}$ Yes See Appendix  B 6.43 6.55
PU Net $4\times 10^{4}$ Yes See Appendix  B 3.69 3.84
Transformer $2\times 10^{4}$ Yes See Appendix  B 3.95 4.18
WaveNet $4\times 10^{4}$ N/A See Appendix  B 4.99 5.16
RNN Seq2Seq $4\times 10^{4}$ Yes See Appendix  B 3.38 3.79
Table 1 shows the denoising scores for all algorithms discussed in Section 3. The case with no denoising is shown in the first row, with its fine denoising score calibrated to 1. We observed that in all cases, the coarse denoising score is slightly higher than the fine denoising score. Based on the results, all traditional algorithms decrease the denoising score because time domain averaging erases high-frequency signals in the region of interest. Meanwhile, all deep learning algorithms efficiently boost the denoising score. Surprisingly, we observed that the FC Net model achieved the best performance with a denoising score of 6.43.
The model training and inference is conducted upon the SDSC Expanse cluster equipped with V100 GPUs. For benchmarking tasks, model training and inference over the denoising score data takes less than O(100 GPU hours). The actual value varies slightly with model complexity and data size.
@@ 4.2 Benchmark 2: dark matter limit
The second benchmark empowers algorithm creators with the capability to conduct a community-standard dark matter search using the science dataset. Following the discussion in the Introduction, the two physics parameters for the dark matter candidate in this paper are the dark matter mass ($m_{a}$) and the dark matter to electromagnetic coupling ($g_{a\gamma\gamma}$). Null results from different dark matter experiments place limits within this parameter space expressed by the shaded regions in Figure 4.2. In the physics community, a better dark matter limit is represented by pushing towards lower values of $g_{a\gamma\gamma}$ at different $m_{a}$. We provide a comprehensive tool necessary for performing the statistical analysis to produce dark matter limits in Figure 4.2. The dark matter mass $m_{a}$ is directly proportional to frequency, therefore the limit-setting procedure is repeated for 11.1 million independent $m_{a}$ from 0.4 neV (100 kHz) to 8 neV (2 MHz). The dark matter limit at each $m_{a}$ is obtained using a frequentist log-likelihood ratio test statistic (TS), with the results depicted as regions in Figure 4.2. The detail of this analysis can be found in Appendix  C.
This analysis is performed twice on the SQUID time series of the science dataset: once without any denoising algorithm which produces the ABRA-TIDMAD Raw limit, and once with FC Net, the top-performing denoising algorithm, which produces the ABRA-TIDMAD Denoised limit. These limits can be directly compared to the previous world-leading ABRA-10cm Run 3 limit, limits obtained by other dark matter experiments, as well as theoretical predictions [25]. While the ABRA-TIDMAD Denoised limit does not outperform the ABRA-10cm Run 3 limit because of hardware and time constraints, it is evident that denoising algorithms significantly improved the dark matter limit by 1-2 orders of magnitude across different $m_{a}$. Although the size of the ABRA-TIDMAD science dataset is only 1% of the ABRA-10cm Run 3 science dataset, the AI denoising algorithm boosted the ABRA-TIDMAD limit to nearly the same level as ABRA-10cm Run 3 and even surpassed the ABRA-10cm Run 3 limits at small $m_{a}$.
Figure 5: Plotted alongside the present, state-of-the-art axion dark matter limits (grey) are the $2\sigma$ exclusion limits for ABRA-10cm Run 3 (pink) [25], ABRA-TIDMAD Raw limit (red), and ABRA-TIDMAD Denoised limit from the trained FC net (maroon). Shaded regions correspond to pairing of dark matter model parameters ($m_{a}$, $g_{a\gamma\gamma}$) that are ruled out by the specified experiment or observation, and the bright yellow region indicates theoretical predictions (See Appendix D for details). The plotting script is modified from the publicly available repository [17].